What is the mass of 22.4 litres of ammonia
1
Expert's answer
2020-10-21T13:19:13-0400

"n=V\/Vm"

n =22,4/22,4=1 mol'

"m=n*M"

M(NH3·H2 О)=35 g /mol'

"m (NH 3 \u00b7 H 2 \u041e) = 1 * 35 = 35 g"

answer: 35g